OpenStudy (anonymous):

an angles measure is 3 times the measure of its compliment. find the measure of the angle and measure of its compliment. can anyone help?

OpenStudy (anonymous):

x+3x=90

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

22.5 will be your answer Jim look at group chat i got a question

OpenStudy (anonymous):

22.5 and 3*22.5

jimthompson5910 (jim_thompson5910):

Let x be the original angle and y be the complement of that angle So y = 3x Since the two angles are complementary, we know that x+y=90, plug in x+3x=90 and solve for x to get 4x=90 x=90/4 x=22.5 So the first angle is 22.5 degrees while the other angle is 67.5 degrees (since 22.5*3 = 67.5)

jimthompson5910 (jim_thompson5910):

which question is that Outkast3r09?

OpenStudy (anonymous):

if you get on group chat i'll ask it's about series tests

OpenStudy (anonymous):

i just don't want to make a question since it really isn't like a problem it's more of